OpenStudy (jenniferkeryakous):

What is the length of a square having the same area as a rectangle that is 24cm by 30cm

OpenStudy (bibby):

a) what is the formula for the area of a square b) what is the area of the rectangle?

OpenStudy (jenniferkeryakous):

I dont know the area of the rectangle or the formula for a square.

OpenStudy (anonymous):

\[side ~of ~\square=\sqrt{30\times24}=?\]

OpenStudy (jenniferkeryakous):

I still dont get what i have to do to get the answer.

OpenStudy (campbell_st):

start by finding the area of the rectangle \[Area = 30 \times 24\] So rectangle area = 720 cm^2 now a square has all sides equal and if the length of a side is l then the area is \[Area = l^2\] now the square and rectangle have the same area, that's started in the question so \[l^2 = 720\] if you take the square root of both sides you can find the side length of the square. hope it helps

OpenStudy (kkutie7):

Area Triangle=1/2*W*H Area of a square= L*W L=W here